OpenStudy (anonymous):

Given the function f(x) = −2x + 10, find x if f(x) = 18.

OpenStudy (anonymous):

\[18 = -2x+10\]First step: add 2x to both sides. What do you get?

OpenStudy (anonymous):

Well, once you get it, the next step is to subtract 18 from both sides.

OpenStudy (anonymous):

i got -6

OpenStudy (anonymous):

Nope. Show me what you got after the first step

OpenStudy (anonymous):

you get 12 if you add 2

OpenStudy (anonymous):

Sorry, add 2x not 2. \[18 = -2x+10\]Now add \(2x\) to both sides.\[18 + 2x=10\]Make sense?

OpenStudy (anonymous):

Now you have \(18+2x=10\) we can subtract 18 to get \(2x=-8\) Hopefully from there it should be ok?

OpenStudy (jhannybean):

\[f(x) = =-2x+10 , f(x)=18\]\[18=-2x+10 \] solving for x, we will eliminte the right side of the equation. \[18-10= 2x+10-10\]\[8=-2x\]\[x=-\frac{ 8 }{ 2 }\] simplify

OpenStudy (anonymous):

-4?

OpenStudy (jhannybean):

Yep.